User Tools

Site Tools


Tid recording to ATNF V011 B1-B7 mounted /data/xraid1

All data have been copied to SWIN V013 B1-B7, to be shipped to Curtin

The ATNF V011 set stays Tid for future experiment recording

03:40:15 On point

04:32:00 LCP 2ndLO frequency corrected

06:39:00 The sky is clear

07:00 End of the schedule

lbaops/lbadec2009/v277ctilog.txt · Last modified: 2015/12/18 16:38 by